The primary winding of the welding transformer has 1000 turns and the secondary winding 110 turns. It is connected to the mains at a voltage of 230 V, giving a secondary current of 120 A. Assume an ideal welding transformer.

a) What is the primary current? How big a fuse should there be behind the socket to which the welding transformer is connected - which of the 10 A, 16 A or 20 A fuses is large enough?

b) What is the secondary voltage?

c) Why is the secondary current in the welding transformer desired high?


1
Expert's answer
2021-09-21T12:02:12-0400

Gives

"N_1=1000\\N_2=110\\V_1=230V"

"I_2=120A"

Part(b)

"\frac{V_1}{V_2}=\frac{N_1}{N_2}\\V_2=\frac{N_2}{N_1}\times V_1"



"V_2=\frac{110}{1000}\times 230"

V2=25.3V

Part(A)

"I_1=\frac{N_2}{N_1}\times I_2"

"I_1=\frac{110}{1000}\times 120=13.2A"

Part(c)

big a fuse should there be behind the socket to which the welding transformer is connected 10 Amp


"I_2=120A" then turn "N_2=110"

High current flow in low turn transformer

"I\alpha \frac{1}{N}"

Current inversely proportional to turns